OpenStudy (anonymous):

Solve the equation x=y+11 2x+y=-2

OpenStudy (anonymous):

211+11

OpenStudy (anonymous):

Soooo, then (0, -11)?

OpenStudy (cj49):

2x+y=-2 x-y=11 (add both the equations) 3x=9 x=3 substitute x=3 in x-y=11

OpenStudy (anonymous):

Am I correct?

OpenStudy (cj49):

??? x=3 to find y ull have to substitute the value of x

OpenStudy (anonymous):

you need to time 2 and 11 theen - 11 and 2

OpenStudy (anonymous):

i mean the sum of 11 +2 x-2

OpenStudy (cj49):

3-y=11 y=11-3 y=9

jhonyy9 (jhonyy9):

@cj49 ATTENTION please 11-3 not is equal 9 !!!!!!!

OpenStudy (cj49):

freaks m sorry i must have been hallucinating...11-3=8